# Web2Code: A Large-scale Webpage-to-Code Dataset and Evaluation Framework for Multimodal LLMs

### v1

@@ Abstract
Multimodal large language models (MLLMs) have shown impressive success across modalities such as image, video, and audio in a variety of understanding and generation tasks. However, current MLLMs are surprisingly poor at understanding webpage screenshots and generating their corresponding HTML code. To address this problem, we propose Web2Code, a benchmark consisting of a new large-scale webpage-to-code dataset for instruction tuning and an evaluation framework for the webpage understanding and HTML code translation abilities of MLLMs. For dataset construction, we leverage pretrained LLMs to enhance existing webpage-to-code datasets as well as generate a diverse pool of new webpages rendered into images. Specifically, the inputs are webpage images and instructions, while the responses are the webpage’s HTML code. We further include diverse natural language QA pairs about the webpage content in the responses to enable a more comprehensive understanding of the web content. To evaluate model performance in these tasks, we develop an evaluation framework for testing MLLMs’ abilities in webpage understanding and web-to-code generation. Extensive experiments show that our proposed dataset is beneficial not only to our proposed tasks but also in the general visual domain, while previous datasets result in worse performance. We hope our work will contribute to the development of general MLLMs suitable for web-based content generation and task automation. Our data and code are available at https://github.com/MBZUAI-LLM/web2code.
@@ 1 Introduction
To demonstrate the utility of our dataset, we train LLaVA-style MLLMs with our dataset included in the instruction finetuning stage. Quantitative results show that finetuning on our dataset not only clearly improves the image-to-HTML-code translation ability of the MLLM, but also leads to slight improvements in the model’s perception and reasoning abilities in general benchmarks. In contrast, including previous datasets without our data conversion leads to degraded general performance, suggesting our dataset is suitable for inclusion in MLLM instruction-tuning for additional capabilities without impacting existing ones.
@@ 2 Related Work
Code Study. There are various code studies related to LLM. Sarker et al. [47] focused on generating code functions through formula hints, trying to improve the robustness of the syntax, and systematically tested the robustness of the syntax. From the perspective of security. Finkman et al. [15] said that these code assistance tools may inadvertently disclose the developer’s proprietary code to the code assistant service provider in the process of helping development, thus they proposed a complementary method to reduce the risk of code leakage while providing effective advice to developers. In addition to code leakage, the code generated by LLM has also caused concerns in industries and other fields. To address this issue, Ye et al. [57] proposed a new zero-shot synthetic code detector based on the similarity between code and its rewritten variants. In the evaluation work on code generation, Du et al. [14] proposed a new computational efficiency benchmark Mercury and a new metric Beyond for the efficiency evaluation of code. They experimentally showed that direct preference optimization can be used as a robust baseline for improving computational efficiency compared to supervised fine-tuning, which paves a promising path for future exploration of efficient code generation.
@@ 3 Dataset Construction
Overview. Our Web2Code instruction tuning dataset construction and instruction generation process involves four key components: (1) Creation of new webpage image-code pair data: We generated high-quality HTML webpage-code pairs following the CodeAlpaca prompt [6] using GPT-3.5 and convert them into instruction-following data. (2) Refinement of existing webpage code generation data: We transform existing datasets including WebSight [22] and Pix2Code [4] into an instruction-following data format similar to LLaVA data [33], so they can be used as instruction-following data to train MLLMs. (3) Creation of a new text question-answer pair data: We generated a new question-answer pair dataset utilizing our new GPT-3.5 generated data from (1) for webpage understanding. (4) Refinement of existing webpage understanding data: We refine the WebSRC [10] question-answer data to improve its quality using the GPT-4. Each component is elaborated in detail as follows:
@@
DWCG: Creation of new webpage image-code pair data for code generation. To augment our dataset with high-quality data, we employed GPT-3.5 to generate 60K HTML pages following the guidelines and prompts in CodeAlpaca [6]. Using Selenium WebDriver, we then created web image screenshots from the generated HTML code. These web image-code pairs were subsequently converted into an instruction-following data format similar to the LLaVA data format [33], enabling their use in training Multimodal Large Language Models (MLLMs). The example of the instruction is shown in Figure 16. The generation of instruction was done in two stages using prompts fed to GPT-4: (a) During stage 1, the prompt shown in Figure 12 resulted in the creation of generic instructions. (b) This was followed by augmenting the instruction from (a) with the GPT generated instructions using the prompt shown in Figure 13 to include stylistic information. This allowed the model to learn two styles: Modern and Bootstrap style as shown in Figure 21 and Figure 22, respectively.
@@
DWCG: Refinement of existing webpage code generation data. To enhance the capability of our model in the task of HTML code generation, we leverage the Pix2code [4] and WebSight [22] datasets. To mitigate the detrimental impact on model performance from random letters in Pix2Code data, we replace these random letters with meaningful text using GPT-4, thereby refining the webpages into diverse webpages encompassing product landing pages, personal portfolios, blogs, and other categories. We then visually rendered each sample by taking screenshots of the browser view of each webpage. Further, we convert all these data into LLaVA instruction following data format using the same strategy as used for DWCG. We note that DWCG and WebSight webpages follow Modern style while Pix2Code follow Bootstrap style.
@@
DWU: Creation of a new question-answer pair data for webpage understanding. For the purpose of fine-tuning our models through an instruction-following manner, we utilized the capabilities of GPT-4 to generate webpage code-based question-answer pairs. We generated 10 question-answer pairs using GPT-4 for a subset of 24.35K webpage data, resulting in a total of 243.5K question-answer data points. This includes, a set of 230K question-answer pairs for GPT-3.5 based webpages, a set of 13.5K newly generated question answer pairs for refined Pix2Code images. These pairs were meticulously crafted to align with our image-based evaluation criteria, ensuring that each question probes specific aspects of the visual and content quality reflected in the generated web images. This strategy enhances the model’s performance by integrating a nuanced understanding of the evaluation parameters into its learning process. Figure 2 shows a qualitative example from our generated question-answer pairs. The prompt used for question-answer generation is shown in Figure 11.
@@
DWU: Refinement of existing webpage understanding data. To increase our instruction-following dataset with high-quality instruction-following examples for webpages, we integrated the WebSRC dataset into our training regime. Prior to inclusion, we meticulously filtered the existing question-and-answer pairs from the WebSRC dataset to ensure relevance and quality. This involves duplication removal and quality optimization as shown in Figure 3. Specifically, we found that WebSRC data contains several questions related to the same answer. To this end, we first removed those duplicates and then employed GPT-4 to assess and enhance the quality of answers. This process not only refined the dataset into 51.5K high-quality instruction data but also ensured that the model’s training was influenced by high-fidelity, instructionally sound data, thereby improving its ability to follow complex web-based instructions.
3.1 Statistics and Analysis
@@ 3.2 Distribution
The 295K question-answer pairs consist of 243.5K GPT-4 based question-answer pairs (DWU Data) and 51.5K question-answer pairs from WebSRC image-based data, as shown in Table 2. Our evaluation dataset comprises 1,198 webpage screenshot images, sourced from diverse origins, including WebSight, Pix2Code, GPT-3.5 based data, and manual processes, ensuring a broad representation of web content. Additionally, we utilize 5,990 "yes" / "no" question-answer pairs generated from the GPT-4 Vision API for our Webpage Understanding Benchmark, as shown in Section 6.
@@ 4.1 Evaluation Metric for HTML Code Generation
WUB: This benchmark comprises 5,990 high-quality question-answer pairs generated from GPT-4 Vision API (See prompt 15), based on 1,198 webpage screenshot images, where each answer is either "yes" or "no". These images are sourced from diverse data origins, including WebSight, Pix2Code, GPT-3.5, and manual processes, ensuring a broad representation of web content. Figure 10 shows a qualitative sample data we used for WUB. We test these pairs on various multimodal image understanding models by comparing the predicted answers to the ground truth, with the final accuracy score serving as the evaluation metric as depicted on the right side of Figure 6. Qualitative data examples in our WUB benchmark are shown in Figure 10.
@@
WCGB: Utilizing the same images as the WUB, this benchmark evaluates a multimodal model tasked with generating HTML code from webpage images based on specific instructions. Unlike traditional code-level evaluations, this benchmark assesses the generated webpage’s fidelity at the image level. We convert the predicted HTML codes back into images using Selenium WebDriver to allow a direct visual comparison with the ground truth images. The evaluation, depicted on the left side of Figure 6, considers 10 different aspects, which are further categorized into four evaluation matrices using the GPT-4 Vision API. This image-level evaluation provides a more accurate measure of the model’s code generation capabilities, acknowledging that identical webpages can be constructed from varying codes. The prompt used for evaluation is shown in Figure 14. This framework consists of 10 distinct criteria, which we group into four categories, each encompassing specific criteria that are scored on a 0-10 scale, as shown in Section D of Appendix.
4.2 Quantitative Evaluation for HTML Code Generation of MLLMs
| LLM Backbone | DWCG | DWU | DWCG | DWU | VSA | CAD | TCC | UII | Overall |
| CystalCoder-7B [37] | - | - | - | - | 3.832 | 3.678 | 3.411 | 3.992 | 3.728 |
| ✓ | - | - | - | 7.812 | 7.899 | 8.138 | 8.112 | 7.990 | |
| ✓ | ✓ | - | - | 8.010 | 8.102 | 8.266 | 8.124 | 8.126 | |
| ✓ | ✓ | ✓ | ✓ | 8.023 | 8.071 | 8.321 | 8.334 | 8.187 | |
| CrystalChat-7B [37] | - | - | - | - | 4.714 | 4.572 | 4.865 | 5.147 | 4.825 |
| ✓ | ✓ | - | - | 7.900 | 8.001 | 8.204 | 8.215 | 8.080 | |
| ✓ | ✓ | ✓ | ✓ | 8.384 | 8.287 | 8.417 | 8.488 | 8.394 | |
| Vicuna1.5-7B [11] | - | - | - | - | 3.042 | 3.250 | 3.333 | 3.167 | 3.198 |
| ✓ | ✓ | ✓ | ✓ | 7.876 | 7.687 | 7.267 | 7.563 | 7.598 | |
| LLaMA3-8B [1] | ✓ | ✓ | ✓ | ✓ | 8.522 | 8.564 | 8.421 | 8.611 | 8.530 |
| LLM Backbone | DWCG | DWU | DWCG | DWU | Accuracy (%) |
| CrystalCoder-7B [37] | - | - | - | - | 73.54 |
| ✓ | - | - | - | 71.81 | |
| ✓ | ✓ | - | - | 73.74 | |
| ✓ | ✓ | ✓ | ✓ | 73.61 | |
| CrystalChat-7B [37] | - | - | - | - | 73.94 |
| ✓ | ✓ | - | - | 73.48 | |
| ✓ | ✓ | ✓ | ✓ | 74.14 | |
| Vicuna1.5-7B [11] | - | - | - | - | 71.12 |
| ✓ | ✓ | ✓ | ✓ | 71.23 | |
| LLaMA3-8B [1] | ✓ | ✓ | ✓ | ✓ | 74.84 |
To be specific, our dataset components have an orthogonal contribution to the overall improvements on both the WUB and the WCGB benchmarks. Table 3 demonstrates improvements in webpage code generation quality when adding DWU, DWCG + DWU, and DWCG + DWU. For example, the results based on instruction-tuned CrystalChat show step-wise improvements on the WCGB benchmark when adding DWCG + DWU, and DWCG + DWU (4.8258.0808.530 on the overall metric). Interestingly, the instruction-tuned CrystalChat-7B with the general domain data only shows poor WCGB performances while it shows on par WUB performance compared to the use of further webpage datasets. Similar trends are also found in the Vicuna1.5-7B backbone, while adding the proposed dataset shows significant improvements, achieving an overall score of 7.598.
Similarly, Table 4 further demonstrates the effectiveness of the proposed dataset on the webpage comprehension capabilities. For example, the four rows of the instruction-tuned CrystalCoder-7B show that not only using DWCG with DWU can prevent degradation in language understanding, but also the full components even enhance the webpage comprehension capabilities. Overall, we found the proposed dataset can enhance both webpage understanding capability and webpage code generation abilities under various LLM backbones, and LLaMA3-8B archives the best performance among all on both webpage code generation and webpage understanding.
4.3 Visualizations for Qualitative Evaluation
As shown in Figure 7, we compare the results between the original image which is the real-world webpage sample, the rendered image generated by using LLM backbones of Vicuna1.5-7B and CrystalChat-7B, respectively. CrystalChat-7B is a code-enhanced LLM and our visualization demonstrates that it achieves the better quality of generation than Vicuna1.5-7B even the performance is slightly worse on general multimodal domain, as presented in Table 6. Moreover, as in Figure 9, our rendered webpage from the model trained on our web dataset closely resembles the original image, indicating the positive impact of the web2code dataset. We further visualize our generation in Figure 9 when the input is a hand-drawn webpage to examine the adaptation ability of our model.
5 General Evaluation of MLLMs Using Web2Code
Setup and Overview. Our model training framework mainly follows the design of LLaVA-1.5 [33] where we leverage the capabilities of both a pre-trained visual encoder, an LLM and a projector to connect visual features into the word embedding space. The model consists of (1) a pre-trained CLIP ViT-L/14 [44] visual encoder with a resolution of 336336 and a patch size of 14, which has good feature representation already aligned with the text embedding space. (2) As for the LLM backbones, we leverage CrystalChat [37] as the base model and compare it with other latest LLM backbones like Vicuna1.5 [11], LLaMA2 [54], LLaMA3 [1] and CrystalCoder [37]111CrystalCoder [37] and CrystalChat [37] are an open-source code LLM pre-trained and instruction-tuned models, respectively. They are trained on publicly available language and code datasets.. Training details and hyperparameters are presented in the Appendix A.
| CrystalChat-7B [37] | - | - | - | - | 1456.53 | 308.21 | 86.86 | 67.77 | 57.84 |
| ✓ | - | - | - | 1438.51 | 292.14 | 87.10 | 68.27 | 58.15 | |
| ✓ | ✓ | - | - | 1478.82 | 297.14 | 86.13 | 67.92 | 57.41 | |
| ✓ | ✓ | ✓ | ✓ | 1449.54 | 279.64 | 86.53 | 68.32 | 57.86 |
Effects of Web2Code on General Domain. Here, we first perform instruction tuning using Web2Code on various LLM backbones and then we evaluate those MLLMs on the general domain of visual language understanding. Throughout extensive experiments under various data configurations, we observed that the proposed dataset Web2Code can be incorporated with the conventional visual language instruction tuning dataset of LLaVA [33] without harming performances on the general domain. Table 5 summarizes the results222We observe that the conventional visual language domain data (i.e., LLaVA) is a crucial component for visual language understanding, i.e., instruction-tuned MLLMs without the general domain data are weak.. Specifically, both proposed Web Understanding data (DWU or DWU) and Web Code Generation data (DWCG or DWCG) do not hurt or even can be beneficial to the visual language understanding. For example, we observed that adding DWU to CrystalChat achieves comparable or even better performances on POPE (86.8687.10), SciQA (67.7768.27), and TextVQA (57.8458.15). Somewhat surprisingly, we further found that adding DWCG can even improve visual language understanding. For example, the second and third rows of CrystalChat show +40.31 and +5.00 points higher improvements in MME-P and MME-C benchmarks, respectively. Moreover, adding refined data DWU and DWCG are still effective in the visual language domain, by achieving comparable (or even better) performances on overall benchmarks. For example, the last row indicates that adding DWU and DWCG preserves comparable performances on overall benchmarks and even achieves +0.4 higher points on the SciQA benchmark.
6 Conclusion
We have presented Web2Code, a benchmark that consists of a high-quality, large-scale webpage-to-code instruction tuning dataset containing 1179.7k entries and an evaluation suite for the webpage understanding and webpage-to-HTML translation abilities of MLLMs. Through extensive experiments, we have demonstrated that our proposed dataset is clearly effective at enhancing these abilities of MLLMs as well as general visual proficiency, while existing datasets lead to inferior performance. We hope our work will attract the community’s attention and facilitate progress toward foundation models serving as virtual assistants for content generation and task automation.
Limitations and Ethics Statement. The Web2Code project provides a comprehensive dataset and evaluation framework for fine-grained multimodal large language models. This can significantly enhance the capabilities of LLMs in understanding and generating web code from instructions, leading to advancements in web development automation, improved coding assistance tools and platforms. By enabling more accurate and context-aware code generation, it can boost productivity for developers and make coding more accessible to beginners. However, the primary limitations of the Web2Code include the potential for biased dataset that may not cover all possible HTML coding scenarios, potentially leading to gaps in model performance, some webpages that include humans may be privacy sensitive. Ensuring high-quality annotations and comprehensive coverage of all possible HTML and code structures is challenging. Also, handling complex, real-world HTML and code scenarios might still be beyond the current capabilities of the models trained on this dataset. Moreover, the proposed evaluation framework may not capture all aspects of the code generation quality, such as code efficiency, readability, or adherence to best practices.
Acknowledgements
We are thankful to Veselin Stoyanov for discussions on multimodal large language models. We also thank the support provided by the MBZUAI IT/corporate service teams (Ian Mathews, John Murphy, Padma Pavani, Tapas Sen, Walid Omari) and CIAI engineering team (Guowei He, Yun Xu, Yue Peng) for organizing High Performance Computing resources and services. Z.S. and E.X. would like to thank the MBZUAI-WIS Joint Program for AI Research. Z.S. also would like to thank the Google Research award grant for the support.
Appendix
Appendix A Training Details and Hyperparameters
Appendix B More Effects of Web2Code on General Domain
| LLM Backbone | MME-P | MME-C | POPE | SciQA | TextVQA |
| CodeLlama-7B [45] | 1345.93 | 258.92 | 85.28 | 61.87 | 55.23 |
| CrystalCoder-7B [37] | 1351.22 | 274.64 | 86.05 | 61.63 | 50.11 |
| CrystalChat-7B [37] | 1478.82 | 297.14 | 86.14 | 67.92 | 57.41 |
| Vicuna1.5-7B [11] | 1488.26 | 268.21 | 87.05 | 69.31 | 56.40 |
| LLaMA2-7B [54] | 1448.58 | 304.29 | 86.87 | 67.87 | 55.62 |
Appendix C Qualitative Data Examples in WUB Benchmark
Appendix D WCGB Criteria
Our proposed WCGB framework consists of 10 distinct criteria, which we group into four categories, each encompassing specific criteria that are scored on a 0-10 scale:

### v2

@@ Abstract
Multimodal large language models (MLLMs) have shown impressive success across modalities such as image, video, and audio in a variety of understanding and generation tasks. However, current MLLMs are surprisingly poor at understanding webpage screenshots and generating their corresponding HTML code. To address this problem, we propose Web2Code, a benchmark consisting of a new large-scale webpage-to-code dataset for instruction tuning and an evaluation framework for the webpage understanding and HTML code translation abilities of MLLMs. For dataset construction, we leverage pretrained LLMs to enhance existing webpage-to-code datasets as well as generate a diverse pool of new webpages rendered into images. Specifically, the inputs are webpage images and instructions, while the responses are the webpage’s HTML code. We further include diverse natural language QA pairs about the webpage content in the responses to enable a more comprehensive understanding of the web content. To evaluate model performance in these tasks, we develop an evaluation framework for testing MLLMs’ abilities in webpage understanding and web-to-code generation. Extensive experiments show that our proposed dataset is beneficial not only to our proposed tasks but also in the general visual domain. We hope our work will contribute to the development of general MLLMs suitable for web-based content generation and task automation. Our data and code are available at https://github.com/MBZUAI-LLM/web2code.
@@ 1 Introduction
To demonstrate the utility of our dataset, we train LLaVA-style MLLMs with our dataset included in the instruction finetuning stage. Quantitative results show that finetuning on our dataset not only clearly improves the image-to-HTML-code translation ability of the MLLM, but also leads to improvements in the model’s perception and reasoning abilities in webpage screenshot understanding as they are closely related. Furthermore, previous datasets like WebSight [22] and Pix2Code [4] present challenges: WebSight includes privacy-sensitive information, while Pix2Code contains random letters in full webpage screenshots. These issues limit MLLMs’ ability to generate coherent text for webpage construction. In contrast, our dataset is more suitable for MLLM instruction fine-tuning, offering enhanced capabilities without compromising existing ones.
@@ 2 Related Work
Code Study. There are various code studies related to LLM. Sarker et al. [47] focuses on generating code functions using formula hints, aiming to enhance syntactic robustness and systematically testing the reliability of the syntax. From the perspective of security, Finkman et al. [15] claims that these code assistance tools may inadvertently disclose the developer’s proprietary code to the code assistant service provider in the process of helping development, thus they propose a complementary method to reduce the risk of code leakage while providing effective advice to developers. In addition to code leakage, the code generated by LLM has also caused concerns in industries and other fields. To address this issue, Ye et al. [57] proposes a new zero-shot synthetic code detector based on the similarity between code and its rewritten variants. In the evaluation work on code generation, Du et al. [14] proposes a new computational efficiency benchmark Mercury and a new metric Beyond for the efficiency evaluation of code. They experimentally show that direct preference optimization can be used as a robust baseline for improving computational efficiency compared to supervised fine-tuning, which paves a promising path for future exploration of efficient code generation.
@@ 3 Dataset Construction
Overview. Our Web2Code instruction tuning dataset construction and instruction generation process involves four key components: (1) Creation of new webpage image-code pair data: We generate high-quality HTML webpage-code pairs following the CodeAlpaca prompt [6] using GPT-3.5 and convert them into instruction-following data. (2) Refinement of existing webpage code generation data: We transform existing datasets including WebSight [22] and Pix2Code [4] into an instruction-following data format similar to LLaVA data [33], so they can be used as instruction-following data to train MLLMs. (3) Creation of a new text question-answer pair data: We generate a new question-answer pair dataset utilizing our new GPT-3.5 generated data from (1) for webpage understanding. (4) Refinement of existing webpage understanding data: We refine the WebSRC [10] question-answer data to improve its quality using the GPT-4. Each component is elaborated in detail as follows:
@@
DWCG: Creation of new webpage image-code pair data for code generation. To augment our dataset with high-quality data, we employ GPT-3.5 to generate 60K HTML pages following the guidelines and prompts in CodeAlpaca [6].111To ensure both quality and diversity in the synthetic data generation, we design a prompting mechanism based on 10 detailed criteria, as detailed in Section 4.1. Using Selenium WebDriver, we then create web image screenshots from the generated HTML code. These web image-code pairs were subsequently converted into an instruction-following data format similar to the LLaVA data format [33], enabling their use in training Multimodal Large Language Models (MLLMs). The example of the instruction is shown in Figure 17. The generation of instruction is done in two stages using prompts fed to GPT-4: (a) During stage 1, the prompt shown in Figure 13 resulted in the creation of generic instructions. (b) This is followed by augmenting the instruction from (a) with the GPT generated instructions using the prompt shown in Figure 14 to include stylistic information. This allows the model to learn two styles: Modern and Bootstrap style as shown in Figure 22 and Figure 23, respectively.
@@
DWCG: Refinement of existing webpage code generation data. To enhance the capability of our model in the task of HTML code generation, we leverage the Pix2code [4] and WebSight [22] datasets. To mitigate the detrimental impact on model performance from random letters in Pix2Code data, we replace these random letters with meaningful text using GPT-4, thereby refining the webpages into diverse webpages encompassing product landing pages, personal portfolios, blogs, and other categories. We then visually render each sample by taking screenshots of the browser view of each webpage. Further, we convert all these data into LLaVA instruction following data format using the same strategy as used for DWCG. We note that DWCG and WebSight webpages follow Modern style while Pix2Code follows Bootstrap style.
@@
DWU: Creation of a new question-answer pair data for webpage understanding. For the purpose of fine-tuning our models in an instruction-following manner, we utilize the capabilities of GPT-4 to generate webpage code-based question-answer pairs. We generate 10 question-answer pairs using GPT-4 for a subset of 24.35K webpage data, resulting in a total of 243.5K question-answer data points. This includes, a set of 230K question-answer pairs for GPT-3.5 based webpages, a set of 13.5K newly generated question answer pairs for refined Pix2Code images. These pairs are meticulously crafted to align with our image-based evaluation criteria, ensuring that each question probes specific aspects of the visual and content quality reflected in the generated web images. This strategy enhances the model’s performance by integrating a nuanced understanding of the evaluation parameters into its learning process. To generate the DWU question-answer pair data, we use only the HTML code shown in Figure 12, which includes the compiled HTML image. Figure 2 shows the compiled HTML image alongside the corresponding question-answer pairs. For training, we input the image and questions together, excluding the HTML code for enhancing webpage understanding capabilities.
@@
DWU: Refinement of existing webpage understanding data. To increase our instruction-following dataset with high-quality instruction-following examples for webpages, we integrate the WebSRC dataset into our training regime. Prior to inclusion, we meticulously filter the existing question-and-answer pairs from the WebSRC dataset to ensure relevance and quality. This involves duplication removal and quality optimization as shown in Figure 3. Specifically, we find that WebSRC data contains several questions related to the same answer. To this end, we first remove those duplicates and then employed GPT-4 to assess and enhance the quality of answers. This process not only refines the dataset into 51.5K high-quality instruction data but also ensures that the model’s training is influenced by high-fidelity, instructionally sound data, thereby improving its ability to follow complex web-based instructions.
@@ 3.2 Distribution
The 295K question-answer pairs consist of 243.5K GPT-4 based question-answer pairs (DWU Data) and 51.5K pairs from WebSRC image-based data, as shown in Table 2. Our evaluation dataset comprises 1,198 webpage screenshot images, sourced from diverse origins, including WebSight, Pix2Code, GPT-3.5-generated data, and manual processes, to ensure a broad representation of web content. Additionally, we utilize 5,990 "yes" / "no" question-answer pairs generated from the GPT-4 Vision API for our Webpage Understanding Benchmark, as in Section 4.1.
@@ 4.1 Evaluation Metric for HTML Code Generation
WUB: This benchmark comprises 5,990 high-quality question-answer pairs generated from GPT-4 Vision API (See prompt 16), based on 1,198 webpage screenshot images, where each answer is either "yes" or "no". These images are sourced from diverse data origins, including WebSight, Pix2Code, GPT-3.5, and manual processes, ensuring a broad representation of web content. Figure 11 shows a qualitative sample data we used for WUB. We test these pairs on various multimodal image understanding models by comparing the predicted answers to the ground truth, with the final accuracy score serving as the evaluation metric as depicted on the right side of Figure 6. Qualitative data examples in our WUB benchmark are shown in Figure 11.
@@
WCGB: Utilizing the same images as the WUB, this benchmark evaluates a multimodal model tasked with generating HTML code from webpage images based on specific instructions. Unlike traditional code-level evaluations, this benchmark assesses the generated webpage’s fidelity at the image level. We convert the predicted HTML codes back into images using Selenium WebDriver to allow a direct visual comparison with the ground truth images. The evaluation, depicted on the left side of Figure 6, considers 10 different aspects, which are further categorized into four evaluation matrices using the GPT-4 Vision API. This image-level evaluation provides a more accurate measure of the model’s code generation capabilities, acknowledging that identical webpages can be constructed from varying codes. The prompt used for evaluation is shown in Figure 15. This framework consists of 10 distinct criteria, which we group into four categories, each encompassing specific criteria that are scored on a 0-10 scale, as follows: